Bob is thinking of a number. Nine more than four times the number is fifteen less than twice the number. What is Bob's number?
Answer by Timnewman(323)

Hi dear,
Bob's number is -12
HERE IS THE WORKING:
Let the number be x,
nine more than four times the number is 9+4x.
Fifteen less than twice the number is 2x-15
If,nine more than four times the numner is fifteen less than twice the number,
Then,
9+4x=2x-15
4x-2x=-15-9
2x=-24
x=-12
Bob's number is -12
HOPE THIS HELPS?
Best wishes,
TIMNEWMAN::